Research Report – November 13th

Science Report 13 November 2019
SOL: 3
Crew 215 – Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler
Science Dome Operations: N/A
Green Hab: Green Hab operations have continued. See Green Hab report.
RAM Operations: N/A
EVA: The third EVA for science saw us returning to the White Moon location at the junction of Brahe Hwy and Cow dung Road (517790E 4254750N) to collect more gypsum samples for ISRU cement experiments. It was noted that, at this location out of sim, at least one vehicle has been driving through the exact locations of our previous sampling efforts. Fortunately, this has no effect on our sampling, but it does suggest that our activities are been monitored extremely closely. See EVA Report.
EVA Suit Maintenance: Nominal. See Operations Report.
Psycho-social Stressors Monitoring: Data collection continues through the hexoskin monitoring vests. In addition, equipment for the retrieval of an injured marstronaut was inventoried prior to a rehearsal for an actual event. Subject to approval, a simulated emergency retrieval of the engineer will be held 14Nov2019 at the conclusion of his engineering EVA.
Additional activities: Nil.

Copyright © The Mars Society. All rights reserved. | Main Site